# revtex4-2Repair the float

### v1

@@ Abstract
We have studied the Kondo effect of a spin-1/2 impurity coupled to a two-dimensional altermagnet host material. To attain the low-temperature many-body Kondo physics of the system, we have performed a numerical renormalization group calculations that allows us to access the spectral properties of the system at zero temperature. The impurity spectral function and the Kondo temperature were calculated for different set of parameters, including Rashba spin-orbit coupling (RSOC) and an external magnetic field. Interestingly, in the RSOC and altermagnetic fields, the hybridization function is spin independent, despite the characteristic broken time-reversal symmetry (TRS) of the altermagnet. This is because the alternating sign of the spin splitting of the bands in the momentum space renders equal contributions for both spin components of the hybridization function. Our results demonstrate that, although the hybridization function is TR symmetric, the Kondo temperature is substantially suppressed by the altermagnet coupling. Moreover, we have investigated the effect of an external magnetic field applied in the altermagnet along different directions. Interestingly, we observe an important restraining of the Kondo peak which depends strongly on the direction of the field. This anisotropic effect is, however, masked if strong Zeeman splitting takes place at the impurity, as it shatter the Kondo-singlet state.
@@ I Introduction
Screening of localized magnetic moments by itinerant electrons in metallic systems is paradigmatic in many-body physics in condensed matter. This phenomenon, known as Kondo effect [1], is crucial for the understanding of both electronic transport and magnetic properties of metallic systems doped with magnetic atoms. The physical mechanism underlying this phenomenon was first proposed by J. Kondo in 1963 [2], as spin-flip scatterings of conduction electrons by magnetic impurities diluted in metallic systems, resulting in a minimal resistivity observed at low temperatures [3]. After more than fifty years of its discovery, Kondo effect has been studied in a variety of systems resulting in an abundance of scientific studies [4, 5, 6, 7].
@@
Quite interestingly, although SOC produces important modification in the energy bands of the host materials, numerical renormalization analysis have shown that, at least in the high density regime, the effect is straightforward. The increase of SOC produce a small widening of the conduction band accompanied by a decreasing of the density of states at the Fermi level. Nonetheless, the effective hybridization between the impurity and the conduction electrons around the Fermi level decreases resulting in a decreasing of $T_{K}$ [14]. It seems, therefore, that despite the fact that SOC produces spin-splitting in the energy bands, because TRS is preserved, it has a minor effect in the Kondo screening [15]. A natural follow up question one may ask is whether this is the case for any system with spin-splitting energy bands, for example, the recently discovered altermagnets [16, 17], with peculiar TRS broken energy bands.
@@
Altermagnets are a newly investigated class of collinear two- and three-dimensional antiferromagnets that display a large broken spin degeneracy, even in the absence of SOC [18, 19, 17]. Since the very first reports on these materials, they are gaining considerable attention in the recent few years. Most of these studies focus on the peculiar properties featured by this emergent class of magnets materials, such as large anisotropic (non-relativistic and) spin-splitting in momentum space with protected symmetry points, accompanied by a zero (net) magnetization. This anisotropic spin-splitting leads to the prediction of remarkable physical properties such as gapless superconducting states with mirage gaps [20], “multi-piezo” effect in ${\rm V}_{2}{\rm Se}_{2}{\rm O}$ monolayer [21], spin-polarized Andreev levels [22], Majorana zero modes with zero net magnetization [23], magnon bands with an alternating chirality splitting [24], among many others exotic properties.
@@
In this paper, we study the Kondo screening of magnetic impurities in altermagnets. More specifically, we aim at investigating how the unusual magnetic properties of these materials modify the Kondo temperature of the system. To this end, we propose a spin-1/2 magnetic impurity coupled to an altermagnet host material described formally by a traditional single-impurity Anderson model (SIAM). To accomplish this, we employ the well-known numerical renormalization group (NRG) [25] that allows us to obtain the relevant low-temperature physical quantities. For instance, within NRG calculations we can obtain the local density of states LDOS at the impurity site, from which we estimate the Kondo temperature of the system. Our results shows that intriguing band structure of the altermagnets has interesting consequences on the Kondo physics: (i) despite the spin-splitted band due to broken TRS, there is no spin splitting in the impurity spectral function. This is because the integration of the self-energy over the entire Brilouin zone renders a spin-independent hybridization function. (ii) By comparing the effects of AM and RSOC to the Kondo temperature, we find that $T_{\rm K}$ is up to two order of magnitude smaller for AM coupling as compared to RSOC, for the same strength. (iii) In the presence of external magnetic field, $\bm{B}$, on the altermagnetic-magnetic impurity system, the suppression on the Kondo peak depends strongly on the direction of the applied field. This anisotropy is akin to what was obtained by one of us in quantum impurity coupled to SOC quantum wires [14].
@@
The rest of this paper is structured as follows. In Sec. II we present the model Hamiltonian and the procedure to calculate the hybridization function. Section. III the NRG results are presented. We summarize and conclude in Sec. IV.
@@ II.1 Hamiltonian model
We consider a hybrid system composed by a spin-$1/2$ impurity coupled to a d-wave AM with RSOC, as illustrated in Fig. 1 (a). The total Hamiltonian of the system can be written in the form of a SIAM as,
@@
where $H_{\rm 0}$, $H_{\rm imp}$ and $H_{\rm hyb}$ represents the host altermagnet material, the magnetic impurity and the coupling between electrons of the altermagnet and the magnetic impurity, respectively. More precisely, $H_{\rm 0}=H_{\rm AM}+H_{\rm RSOC}+H_{\rm Z}$, where $H_{\rm AM}$ describes, for instance, a d-wave altermagnet on a substrate [19, 17, 26]. $H_{\rm RSOC}$ accounts for the contribution of the Rashba SOC (RSOC) induced by an inversion asymmetry, which can naturally arise due to the substrate on which the AM is deposited. Finally, $H_{\rm Z}$ represents an external magnetic field. More explicitly, each of these terms are given by
$H_{\rm AM}=2t(\cos k_{x}\!+\!\cos k_{y})\sigma_{0}\!+\!2t_{AM}(\cos k_{x}\!-\!\cos k_{y})\sigma_{z},$
@@
Here, $t$ the hopping amplitude parameter and $t_{AM}$ stands for the AM exchange interaction. We have written $H_{\rm RSOC}$ as an effective $k$-depended magnetic field ${\bm{B}}_{\rm SOC}({\bm{k}})=2\lambda(\sin k_{y},-\sin k_{x})$ is induced by the Rashba field, with $\lambda$ standing for the strength of the interaction. The last term of Eq. (II.1), $H_{Z}$, describes the Zeeman effect in the AM material due to an external magnetic field ${\bm{B}}$, in which $g$ is the effective $g$-factor of the AM. In all these terms, $\sigma_{i}$ represents the Pauli matrices, including the identity $\sigma_{0}$ and $\mathbf{S}=\hbar\bm{\sigma}/2$.
The second term of the Eq. (1) is given by,
@@
where, $\varepsilon_{d}$ is the impurity on-site energy for spin $\sigma$ ($\uparrow$ or $\downarrow$), $g_{\rm imp}$ is the impurity $g$-factor, $U$ represents the $e$-$e$ interaction at the impurity site, and $n_{\sigma}=d^{\dagger}_{\sigma}d_{\sigma}$ is the electron number operator with spin $\sigma$.
@@
$H_{\rm hyb}\!=\!V\sum_{k\sigma}(d^{\dagger}_{\sigma}c_{k\sigma}+H.c.).$ (4)
@@ II.2 Energy band properties
The main features of the electronic bands given by Eq. II.1 is shown in Fig. 1(b)-1(e), for different sets of parameter configurations. Fig. 1(b) shows the pristine case, which shows the alternating spin-splitting along the momentum and with degenerate points at $\Gamma$ and also at the four corners of the Brillouin zone. In Fig. 1 (c) we show the case where $\lambda\neq$ 0 (while ${\bm{B}}$=0), which fully breaks the spin degeneracy of the bands, except at $\Gamma$ points and at the four corners. In Fig. 1(d)-1(e) a ${\bm{B}}$ is considered along different directions (with $\lambda$=0), where now there is no degeneracy at $\Gamma$ and at the boundaries of the Brillouin zone. Below, we will return to the host band structure with a deeper analysis under different interactions.
@@
Figure 2: Energy bands along the $M$-$\Gamma$-$M^{\prime}$ path for different parameters configurations: (a) ${\bm{B}}=0$, (b) ${\bm{B}}=0.05\bm{\hat{\textbf{x}}}$, (c) ${\bm{B}}=0.05\bm{\hat{\textbf{y}}}$ and (d) ${\bm{B}}=0.05\bm{\hat{\textbf{z}}}$. For all panels, $t$=0.5, $t_{AM}$=0.1, $\lambda$=0.2, a Néel vector along the $z$-direction and the color map is associated to the spin-projection, $\langle S_{i}\rangle$, for each band. The insets correspond to each respective cases with $\lambda$=0.0.
@@
Let us further analyze the host band structure in the presence of the external magnetic field. For the sake of numerical convenience, throughout the paper we will set, $g=k_{B}=1.0=\mu_{B}=\hbar=1.0$. Fig. 2(a)-2(d) show the electronic band structure along the path $M$-$\Gamma$-$M^{\prime}$, which is a good representative of the symmetries of the bands. In the absence of magnetic field and RSOC, the system shows an intriguing broken TRS with a momentum dependent alternating spin-splitting, as shown in Fig. 1(b). The band degeneracy along the path $M$-$\Gamma$-$M^{\prime}$ is broken when RSOC is considered. However, they are still degenerated at the $\Gamma$ and $M$ ($M^{\prime}$) points, which is a consequence of the $C_{4z}\mathcal{T}$ symmetry that still holds [26]. This feature can be clearly seen in Fig. 2 (a), where we show the band structure for ${\bm{B}}$=0 and $\lambda=0.2$ ($\lambda=0.0$ for the inset). Moreover, the spin projection $\langle S_{z}\rangle$ is zero except at the spin-degenerate points, where the value is exactly $\langle S_{z}\rangle$=0.5 (and -0.5). Indeed, the spin projection would be non-zero (and equal, $\langle S_{x}\rangle$=$\langle S_{y}\rangle$) only along the plane ($x-y$) (not shown in the panel). For an applied magnetic field in the plane $x$-$y$, say, at the $x$-(or $y$-direction), $C_{4z}\mathcal{T}$-symmetry is broken as shown in Fig. 2(b)-2(c). We observe clear spin splittings of the bands at all points including $\Gamma$ and $M$ ($M^{\prime}$) which acquires a finite gap of $2g|{\bm{B}}|$, even for the $\lambda$=0 case (show in the insets). However, differently from the $\lambda$=0, the presence of RSOC induces a spin mixing of the bands with a broken inversion symmetry along $M$-$\Gamma$ and $\Gamma$-$M^{\prime}$ path. Moreover, by alternating the direction of the in-plane ${\bm{B}}$ field, the band structure around $\Gamma$ point is reversed, with significant changes in the spin projection along the direction of the applied ${\bm{B}}$ for $\lambda\neq 0$, with spin projections inverted at the $\Gamma$ an $M$ ($M^{\prime}$) points. At last, in Fig. 2(d) we show the case when $\bm{B}$ is applied along the $z$-direction. As previously observed, the system is TRS-broken with a clear spin splitting. However, differently from the in-plane case, the system still has a symmetry along the cut line $M$-$\Gamma$-$M$. From these results, we can confirm that the interplay between RSOC and magnetic field introduces important modifications to the altermagnet electronic bands, with a relevant anisotropy that can be crucial for the Kondo physics, as it will be further explored in the following sections.
@@ II.3 Hybridization Function
${\bm{G}}_{\rm imp}(\omega)\!=\!\left[(\varepsilon_{d}-\omega)\sigma_{0}-\bm{\Sigma}^{(0)}(\omega)-\bm{\Sigma}^{({\rm int})}(\omega)\right]^{-1}.$ (5)
Here, $\bm{\Sigma}^{(0)}(\omega)=\sum_{k}\bm{V}\bm{G}_{\rm host}({\bm{k}},\omega)\bm{V}^{\dagger}$ represents the non-interacting self-energy [28], with $\hat{V}=V\sigma_{0}$ and $\bm{G}_{\rm host}(k,\omega)=\left[\omega\sigma_{0}-H_{0}\right]^{-1}$, while $\bm{\Sigma}^{(\rm int)}(\omega)$ is the interacting part (or proper) self-energy. With the non-interacting self-energy, the hybridization function can be written as,
@@
where $\eta\rightarrow 0$. The integration defined in Eq. (6) is highly peaked for small values of $\omega$, therefore a high precision numerical integration is desirable to avoid numerical flaws. To achieve this purpose, we have used cubature rules as implemented in the Cuba package [29]. Note that $\bm{\Gamma}(\omega)$ is also a $2\times 2$ matrix and, in the presence of an external in-plane magnetic field, there is a spin-mixing rendering non-diagonal elements. To deal with the spin-mixing channels, a non-trivial Wilsonian RG chain is required [14]. As the hybridization matrix has, in general, non-zero off-diagonal complex elements, it is useful to decompose it in terms of Pauli matrices as, $\bm{\Gamma}(\omega)=\sum_{i}d_{i}(\omega)\sigma_{i}$ ($i$=0, x, y, z). For convenience, we hereafter will rescale the bandwidth to unity ($D$=1), such that the relevant energy range is defined as $-1<\omega<1$, a standard procedure in impurity solvers. We also assume the impurity-altermagnet coupling to be $V=$0.1.
@@
Figure 3: Hybridization function $\hat{\Gamma}(\omega)$: (a) For ${\lambda}$=0.0 and $t_{AM}\neq$ 0; (b) $t_{AM}$=0.0 and $\lambda\neq$0. For both panels $\bm{B}$=0, and notice that we are showing the trace of $d_{0}\sigma_{0}$, as the diagonal terms are equal and there is no off-diagonal terms.
@@
In Fig. 3 (a) and (b), we show $\Gamma(\omega)$ for different values of $t_{AM}$ and $\lambda$, respectively, in absence of $\bm{B}$. Notice that regardless of the value of $\lambda$ or $t_{AM}$, the diagonal terms for both panels are equal, as there is no spin-mixing terms in the hybridization matrix. Therefore, we only showed the trace of $d_{0}(\omega)\sigma_{0}\propto 2d_{0}$, which is proportional to the total density of states. When $t_{AM}$ is considered (with $\lambda$=0), one can observe a strong suppression of the peak around $\omega=0$, with the development of a splitting. This suppression will be responsible to signatures in the Kondo temperature, as it will be discussed in the following section. For the case when RSOC is non-zero (with $t_{AM}$=0), shown in Fig. 3 (b), we can also notice a peak suppression similar to the Fig. 3 (a), however, the suppression is less pronounced, and there are peaks at the edge associated to the van Hove singularities. Since they are far away from from the Fermi level, these van Hove singularities at the edges near the edge of conduction band is unimportant for the Kondo screening. The more pronounced suppression of the hybridization function near the Fermi level produced by $t_{\rm AM}$ anticipates the stronger suppression of Kondo screening caused by the altermagnet as compared to RSOC.
@@
Figure 4: Decomposition of the matrix hybridization function $\hat{\Gamma}(\omega)$ along different Pauli matrices $\sigma_{i}$ with corresponding weights $d_{i}$. (a) For ${\bm{B}}$=0, (b) for a magnetic field along $\bm{\hat{\textbf{x}}}$, (c) for a magnetic field along $\bm{\hat{\textbf{y}}}$ and (c) for a magnetic field along $\bm{\hat{\textbf{z}}}$. The solid line is for $\lambda$=0, while the dashed line stands for $\lambda$=0.2. For panels (b)-(d), we have assumed $t_{\rm AM}=0.1$ and $B=0.02$.
@@
Now, we set the value of $t_{AM}$=0.1 and investigate how the presence of $\bm{B}$ affects the hybridization matrix. In Fig. 4, we show $\bm{\Gamma}(\omega)$ decomposed for different parameter configuration of ${\bm{B}}$ and $\lambda$. In the absence of a magnetic field and RSOC ($\lambda$=0), $d_{0}$ is the only nonzero value as shown by the black solid line in Fig. 4 (a). However, notice that this result holds even for the case of $\lambda$=0.2 (dashed line). This behavior is because (i) the impurity is spin-independently coupled to the host, (ii) the AM Hamiltonian [$H_{0}$ in Eq. (1)] preserves $C_{4z}\mathcal{T}$ symmetry and SOC is time-reversal invariant, rendering thus a diagonal and spin-independent $\bm{\Gamma}(\omega)$. Nonetheless, notice that because $\lambda\neq$0, we can still see the emergent peaks close to the edges of the conduction band associated to the appearance of van Hove singularities in the conduction band around the $M$-point. For an in-plane magnetic field Fig. 4 (b)-(c), the off-diagonal elements of $\hat{\Gamma}(\omega)$ will contribute to non-zero values for the coefficients $d_{y}$ ($d_{x}$) for a magnetic field applied along $x$ ($y$) conversely, while there is also a finite value for $d_{z}$ that changes sign, i.e. $d_{z}\rightarrow-d_{z}$. Interestingly, for $\bm{B}\parallel\bm{\hat{\textbf{z}}}$ shown in Fig. 4 (d), the only nonzero components are along $d_{0}$ and $d_{z}$, even when $\lambda\neq$0 (which intertwine the spin channels but does not break TRS). Furthermore, notice that because of the interplay between the AM exchange interaction and the Zeeman field, there is a enhancement of the $d_{z}$ component, and also a marked double splitting in the peaks of $d_{0}$, which will strongly reflect in the impurity spectral function.

### v2

@@ Abstract
We have studied the Kondo effect of a spin-1/2 impurity coupled to a two-dimensional altermagnet host material. To attain the low-temperature many-body Kondo physics of the system, we have performed a numerical renormalization group calculations that allows us to access the spectral properties of the system at zero temperature. The impurity spectral function and the Kondo temperature were calculated for different set of parameters, including Rashba spin-orbit coupling (RSOC) and an external magnetic field. Interestingly, in the RSOC and altermagnetic fields, the hybridization function is spin independent, despite the characteristic broken time-reversal symmetry of the altermagnet. This is because the alternating sign of the spin splitting of the bands in the momentum space renders equal contributions for both spin components of the hybridization function. Our results demonstrate that, although the hybridization function is time-reversal symmetric, the Kondo temperature is substantially suppressed by the altermagnet coupling. Moreover, we have investigated the effect of an external magnetic field applied in the altermagnet along different directions. Interestingly, we observe an important restraining of the Kondo peak which depends strongly on the direction of the field. This anisotropic effect is, however, masked if strong Zeeman splitting takes place at the impurity, as it shatters the Kondo-singlet state.
@@ I Introduction
Screening of localized magnetic moments by itinerant electrons in metallic systems is paradigmatic in many-body physics in condensed matter. This phenomenon, known as the Kondo effect [1], is crucial for understanding both electronic transport and magnetic properties of metallic systems doped with magnetic atoms. The physical mechanism underlying this phenomenon was first proposed by Kondo in 1963 [2], as spin-flip scatterings of conduction electrons by magnetic impurities diluted in metallic systems, resulting in a minimal resistivity observed at low temperatures [3]. After more than fifty years of its discovery, Kondo effect has been studied in a variety of systems resulting in an abundance of scientific studies [4, 5, 6, 7].
@@
Quite interestingly, although SOC produces an important modification in the energy bands of the host materials, numerical renormalization analysis has shown that, at least in the high density regime, the effect is straightforward. The increase of SOC produces a small widening of the conduction band accompanied by a decrease of the density of states at the Fermi level. Nonetheless, the effective hybridization between the impurity and the conduction electrons around the Fermi level decreases resulting in a decreasing of $T_{K}$ [14]. It seems, therefore, that despite the fact that SOC produces spin-splitting in the energy bands, because TRS is preserved, it has a minor effect in the Kondo screening [15]. A natural follow up question one may ask is whether this is the case for any system with spin-splitting energy bands, for example, the recently discovered altermagnets [16, 17, 18, 19], with peculiar TRS broken energy bands.
@@
Altermagnets are a class of collinear two- and three-dimensional antiferromagnets [20] that display a large broken spin degeneracy, even in the absence of SOC [21, 22, 19], and they have gained considerable attention recently. Most of these studies have focused on the peculiar properties featured by this emergent class of magnets materials, such as large anisotropic (non-relativistic) and spin-splitting in momentum space with protected symmetry points, accompanied by a zero (net) magnetization. This anisotropic spin-splitting leads to the prediction of remarkable physical properties such as gapless superconducting states with mirage gaps [23], “multi-piezo” effect in ${\rm V}_{2}{\rm Se}_{2}{\rm O}$ monolayer [24], spin-polarized Andreev levels [25], Majorana zero modes with zero net magnetization [26], magnon bands with an alternating chirality splitting [27], pronounced thermal transport [28], among many others exotic properties.
@@
In this paper, we study the Kondo screening of magnetic impurities in altermagnets. More specifically, we aim at investigating how the unusual magnetic properties of these materials modify the Kondo temperature of the system. To this end, we propose a spin-1/2 magnetic impurity coupled to an altermagnet host material described formally by a traditional single-impurity Anderson model (SIAM). To accomplish this, we employ the well-known numerical renormalization group (NRG) [29] that allows us to obtain the relevant low-temperature physical quantities. For instance, within NRG calculations we can obtain the local density of states LDOS at the impurity site, from which we estimate the Kondo temperature of the system. Our results shows that the intriguing band structure of the altermagnets has interesting consequences on the Kondo physics: (i) despite the spin-splitted band due to broken TRS, there is no spin splitting in the impurity spectral function. This is because the integration of the self-energy over the entire Brillouin zone renders a spin-independent hybridization function. (ii) By comparing the effects of AM and RSOC to the Kondo temperature, we find that $T_{\rm K}$ is up to two order of magnitude smaller for AM coupling as compared to RSOC, for the same strength. (iii) In the presence of external magnetic field, $\bm{B}$, on the altermagnetic-magnetic impurity system, the suppression on the Kondo peak depends strongly on the direction of the applied field. This anisotropy is akin to what was obtained by one of us in quantum impurity coupled to SOC quantum wires [14].
@@
The rest of this paper is structured as follows. In Sec. II we present the model Hamiltonian and the procedure to calculate the hybridization function. Sec. III the NRG results are presented. We summarize and conclude in Sec. IV.
@@ II.1 Hamiltonian model
We consider a hybrid system composed by a spin-$1/2$ impurity coupled to a $d$-wave AM with RSOC [30], as illustrated in Fig. 1 (a). The total Hamiltonian of the system can be written in the form of a SIAM as,
@@
where $H_{\rm 0}$, $H_{\rm imp}$ and $H_{\rm hyb}$ represent the host altermagnet material, the magnetic impurity and the coupling between electrons of the altermagnet and the magnetic impurity, respectively. More precisely, $H_{\rm 0}=H_{\rm AM}+H_{\rm RSOC}+H_{\rm Z}$, where $H_{\rm AM}$ describes, for instance, a $d$-wave altermagnet on a substrate [22, 19, 30]. $H_{\rm RSOC}$ accounts for the contribution of the Rashba SOC (RSOC) induced by an inversion asymmetry, which can naturally arise due to the substrate on which the AM is deposited. Finally, $H_{\rm Z}$ represents an external magnetic field. More explicitly, each of these terms are given by
$H_{\rm AM}=2t(\cos k_{x}\!+\!\cos k_{y})\sigma_{0}\!+\!2t_{\rm AM}(\cos k_{x}\!-\!\cos k_{y})\sigma_{z},$
@@
Here, $t$ is the hopping amplitude parameter and $t_{\rm AM}$ stands for the AM exchange interaction. We have written $H_{\rm RSOC}$ as an effective $k$-dependent magnetic field ${\bm{B}}_{\rm SOC}({\bm{k}})=2\lambda(\sin k_{y},-\sin k_{x})$ induced by the Rashba field, with $\lambda$ standing for the strength of the interaction. The last term of Eq. (II.1), $H_{Z}$, describes the Zeeman effect in the AM material due to an external magnetic field ${\bm{B}}$, in which $g$ is the effective $g$-factor of the AM. In all these terms, $\sigma_{i}$ represents the Pauli matrices, including the identity $\sigma_{0}$ and $\mathbf{S}=\hbar\bm{\sigma}/2$.
The second term of Eq. (1) is given by
@@
where, $\varepsilon_{d}$ is the impurity on-site energy for spin $\sigma$ ($\uparrow$ or $\downarrow$), $g_{\rm imp}$ is the impurity $g$ factor, $U$ represents the $e$-$e$ interaction at the impurity site, and $n_{\sigma}=d^{\dagger}_{\sigma}d_{\sigma}$ is the electron number operator with spin $\sigma$.
@@
$H_{\rm hyb}\!=\!V\sum_{k\sigma}(d^{\dagger}_{\sigma}c_{k\sigma}+H.c.),$ (4)
@@ II.2 Energy band properties
The main features of the electronic bands given by Eq. II.1 are shown in Figs. 1(b)-1(e), for different sets of parameter configurations. Fig. 1(b) shows the pristine case, which shows the alternating spin-splitting along the momentum and with degenerate points at $\Gamma$ and also at the four corners of the Brillouin zone. In Fig. 1 (c), we show the case where $\lambda\neq$ 0 (while ${\bm{B}}$=0), which fully breaks the spin degeneracy of the bands, except at $\Gamma$ points and at the four corners. In Figs. 1(d)-1(e), a ${\bm{B}}$ is considered along different directions (with $\lambda$=0), where now there is no degeneracy at $\Gamma$ and at the boundaries of the Brillouin zone. Below, we will return to the host band structure with a deeper analysis under different interactions.
@@
Figure 2: Energy bands along the $M$-$\Gamma$-$M^{\prime}$ path for different parameters configurations: (a) ${\bm{B}}=0$, (b) ${\bm{B}}=0.05\bm{\hat{\textbf{x}}}$, (c) ${\bm{B}}=0.05\bm{\hat{\textbf{y}}}$, and (d) ${\bm{B}}=0.05\bm{\hat{\textbf{z}}}$. For all panels, $t$=0.5, $t_{\rm AM}$=0.1, $\lambda$=0.2, a Néel vector along the $z$-direction and the color map is associated to the spin-projection, $\langle S_{i}\rangle$, for each band. The insets correspond to each respective cases with $\lambda$=0.0.
@@
Let us further analyze the host band structure in the presence of the external magnetic field. For the sake of numerical convenience, throughout the paper we will set, $g=k_{B}=1.0=\mu_{B}=\hbar=1.0$. Fig. 2(a)-2(d) show the electronic band structure along the path $M$-$\Gamma$-$M^{\prime}$, which is a good representative of the symmetries of the bands. In the absence of magnetic field and RSOC, the system shows an intriguing broken TRS with a momentum dependent alternating spin-splitting, as shown in Fig. 1(b). The band degeneracy along the path $M$-$\Gamma$-$M^{\prime}$ is broken when RSOC is considered. However, they are still degenerated at the $\Gamma$ and $M$ ($M^{\prime}$) points, which is a consequence of the $C_{4z}\mathcal{T}$ symmetry that still holds [30]. This feature can be clearly seen in Fig. 2 (a), where we show the band structure for ${\bm{B}}$=0 and $\lambda=0.2$ ($\lambda=0.0$ for the inset). Moreover, the spin projection $\langle S_{z}\rangle$ is zero except at the spin-degenerate points, where the value is exactly $\langle S_{z}\rangle$=0.5 (and -0.5). Indeed, the spin projection would be non-zero (and equal, $\langle S_{x}\rangle$=$\langle S_{y}\rangle$) only along the plane ($x-y$) (not shown in the panel). For an applied magnetic field in the plane $x$-$y$, say, at the $x$-(or $y$-direction), $C_{4z}\mathcal{T}$-symmetry is broken as shown in Fig. 2(b) and 2(c). We observe clear spin splittings of the bands at all points including $\Gamma$ and $M$ ($M^{\prime}$), which acquires a finite gap of $2g|{\bm{B}}|$, even for the $\lambda$=0 case (shown in the insets). However, differently from the $\lambda$=0, the presence of RSOC induces a spin mixing of the bands with a broken inversion symmetry along the $M$-$\Gamma$ and $\Gamma$-$M^{\prime}$ path. Moreover, by alternating the direction of the in-plane ${\bm{B}}$ field, the band structure around the $\Gamma$ point is reversed, with significant changes in the spin projection along the direction of the applied ${\bm{B}}$ for $\lambda\neq 0$, with spin projections inverted at the $\Gamma$ an $M$ ($M^{\prime}$) points. Lastly, in Fig. 2(d) we show the case when $\bm{B}$ is applied along the $z$-direction. As previously observed, the system is TRS-broken with a clear spin splitting. However, differently from the in-plane case, the system still has a symmetry along the cut line $M$-$\Gamma$-$M$. From these results, we can confirm that the interplay between RSOC and magnetic field introduces important modifications to the altermagnet electronic bands, with a relevant anisotropy that can be crucial for the Kondo physics, as will be further explored in the following sections.
@@ II.3 Hybridization function
${\cal G}_{\rm imp}(\omega)\!=\!\left[(\varepsilon_{d}-\omega)\sigma_{0}-\bm{\Sigma}^{(0)}(\omega)-\bm{\Sigma}^{({\rm int})}(\omega)\right]^{-1}.$ (5)
Here, $\bm{\Sigma}^{(0)}(\omega)=\sum_{k}\bm{V}\bm{G}_{\rm host}({\bm{k}},\omega)\bm{V}^{\dagger}$ represents the non-interacting self-energy [32], with $\hat{V}=V\sigma_{0}$ and $\bm{G}_{\rm host}(k,\omega)=\left[\omega\sigma_{0}-H_{0}\right]^{-1}$, while $\bm{\Sigma}^{(\rm int)}(\omega)$ is the interacting part (or proper) self-energy, which is obtained within the NRG calculation. With the non-interacting self-energy, the hybridization function can be written as,
@@
where $\eta\rightarrow 0$. The integration defined in Eq. (6) is highly peaked for small values of $\omega$, therefore a high precision numerical integration is desirable to avoid numerical flaws. To achieve this purpose, we have used cubature rules as implemented in the Cuba package [33]. Note that $\bm{\Gamma}(\omega)$ is also a $2\times 2$ matrix and, in the presence of an external in-plane magnetic field, there is a spin-mixing rendering non-diagonal elements. To deal with the spin-mixing channels, a non-trivial Wilsonian RG chain is required [14]. As the hybridization matrix has, in general, non-zero off-diagonal complex elements, it is useful to decompose it in terms of Pauli matrices as, $\bm{\Gamma}(\omega)=\sum_{i}d_{i}(\omega)\sigma_{i}$ ($i$=0, x, y, z). For convenience, we hereafter will rescale the bandwidth to unity ($D$=1), such that the relevant energy range is defined as $-1<\omega<1$, a standard procedure in quantum impurity solvers. We also assume the impurity-altermagnet coupling to be $V=$0.1.
@@
Figure 3: Hybridization function $\hat{\Gamma}(\omega)$: (a) for ${\lambda}$=0.0 and $t_{\rm AM}\neq$ 0; (b) $t_{\rm AM}$=0.0 and $\lambda\neq$0. For both panels $\bm{B}$=0, and notice that we are showing the trace of $d_{0}\sigma_{0}$, as the diagonal terms are equal and there are no off-diagonal terms.
@@
In Figs. 3 (a) and 3(b), we show $\Gamma(\omega)$ for different values of $t_{\rm AM}$ and $\lambda$, respectively, in the absence of $\bm{B}$. Notice that regardless of the value of $\lambda$ or $t_{\rm AM}$, the diagonal terms for both panels are equal, as there is no spin-mixing terms in the hybridization matrix. Therefore, we only showed the trace of $d_{0}(\omega)\sigma_{0}\propto 2d_{0}$, which is proportional to the total density of states. When $t_{\rm AM}$ is considered (with $\lambda$=0), one can observe a strong suppression of the peak around $\omega=0$, with the development of a splitting. This suppression will be responsible to signatures in the Kondo temperature, as it will be discussed in the following section. For the case when RSOC is non-zero (with $t_{\rm AM}$=0), shown in Fig. 3 (b), we can also notice a peak suppression similar to the Fig. 3 (a), however, the suppression is less pronounced, and there are peaks at the edge associated to the van Hove singularities. Since they are far away from the Fermi level, these van Hove singularities at the edges near the edge of the conduction band are unimportant for the Kondo screening. The more pronounced suppression of the hybridization function near the Fermi level produced by $t_{\rm AM}$ anticipates the stronger suppression of Kondo screening caused by the altermagnet as compared to RSOC.
@@
Figure 4: Decomposition of the matrix hybridization function $\hat{\Gamma}(\omega)$ along different Pauli matrices $\sigma_{i}$ with corresponding weights $d_{i}$. (a) For ${\bm{B}}$=0, (b) for a magnetic field along $\bm{\hat{\textbf{x}}}$, (c) for a magnetic field along $\bm{\hat{\textbf{y}}}$ and (d) for a magnetic field along $\bm{\hat{\textbf{z}}}$. The solid line is for $\lambda$=0, while the dashed line stands for $\lambda$=0.2. For (b)-(d), we have assumed $t_{\rm AM}=0.1$ and $B=0.02$.
@@
Now we set the value of $t_{\rm AM}$=0.1 and investigate how the presence of $\bm{B}$ affects the hybridization matrix. In Fig. 4, we show $\bm{\Gamma}(\omega)$ decomposed for different parameter configurations of ${\bm{B}}$ and $\lambda$. In the absence of a magnetic field and RSOC ($\lambda$=0), $d_{0}$ is the only nonzero value, as shown by the black solid line in Fig. 4 (a). However, notice that this result holds even for the case of $\lambda$=0.2 (dashed line). This behavior is because (i) the impurity is spin-independently coupled to the host and (ii) the AM Hamiltonian [$H_{0}$ in Eq. (1)] preserves $C_{4z}\mathcal{T}$ symmetry and SOC is time-reversal invariant, thus rendering a diagonal and spin-independent $\bm{\Gamma}(\omega)$. Nonetheless, notice that because $\lambda\neq$0, we can still see the emergent peaks close to the edges of the conduction band associated to the appearance of van Hove singularities in the conduction band around the $M$-point. For an in-plane magnetic field Fig. 4 (b) and 4(c), the off-diagonal elements of $\hat{\Gamma}(\omega)$ will contribute to non-zero values for the coefficients $d_{y}$ ($d_{x}$) for a magnetic field applied along $x$ ($y$) conversely, while there is also a finite value for $d_{z}$ that changes signs, i.e., $d_{z}\rightarrow-d_{z}$. Interestingly, for $\bm{B}\parallel\bm{\hat{\textbf{z}}}$ shown in Fig. 4 (d), the only nonzero components are along $d_{0}$ and $d_{z}$, even when $\lambda\neq$0 (which intertwine the spin channels but does not break TRS). Furthermore, notice that because of the interplay between the AM exchange interaction and the Zeeman field, there is an enhancement of the $d_{z}$ component, and also a marked double splitting in the peaks of $d_{0}$, which will strongly reflect in the impurity spectral function.